OpenStudy (johan14th):

Can someone help me find the integration by substitution of a problem, I'm having a bit of problems with the trig in there. Thanks :)

OpenStudy (johan14th):

\[\int\limits_{0}^{\frac{ \pi }{ 4 }}\sin (2x)dx=\]

OpenStudy (johan14th):

so far i have u=2x du=2dx \[\frac{ 1 }{ 2 }\int\limits_{0}^{\frac{ \pi}{ 4 }}\sin (u)2du\] \[-\frac{ 1 }{ 2 }\cos (u)+c\] \[-\frac{ 1 }{ 2 }\cos (2x)\] F(b)-F(a) \[-\frac{ 1 }{ 2 }\cos(2(\frac{ \pi }{ 4 }))-[-\frac{ 1 }{ 2 }\cos (2(0))] \]

OpenStudy (johan14th):

I'm having a bit of problem with the next part